OpenStudy (anonymous):

fan and metal please help 1st person to get it right What is the slope of the line passing through the points (–1, 3) and (4, –7)? A.2 B.3/4 C.-4/3 D.-2

OpenStudy (mathmale):

Please do NOT introduce your posts with "fan and metal please help 1st person to get it right." Most importantly, you are not allowed on OpenStudy to ask for answers, and other users are not allowed to give them to you. To solve this problem, look at the 2 given points. Find the slope of the line that connects these points. The slope will be denoted by m. You want to find the equation y = mx + b. How would you do that, using one of the given points for your data (x,y)?

OpenStudy (anonymous):

ok i wont any more

OpenStudy (anonymous):

I am not good at thoughs

OpenStudy (mathmale):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? The appropriate formula is

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (mathmale):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

K thanks alex

OpenStudy (anonymous):

its -2

OpenStudy (mathmale):

It doesn't matter which point is the first, or which is the second. You could let the point (-1,3) be the first, so that x1=-1 and y1=3. Please do the same thing for the second given point. What is the slope of the line connecting the 2 given points?

OpenStudy (mathmale):

Mind explaining how you decided that "it's -2"?

OpenStudy (anonymous):

I did -7 - 3 / 4- (-1) = -10/5 =-2